OpenStudy (anonymous):

A shop sells compact disks for $11 per piece. If each disk costs $5 to the shop owner how much profit percent does the shop owner earn?

OpenStudy (anonymous):

(6/11)(100)

OpenStudy (anonymous):

revenue - cost = profit 11 - 5 = 6 (profit/revenue)*100 = (6/11)*100 = 54.54%

OpenStudy (anonymous):

good I got the same answer but somehow I came up with it different

OpenStudy (anonymous):

Glad to hear it. If you want to solve problems faster during the SAT test math sections, try our SAT app for your calculator: http://download.cnet.com/SAT-Operating-System-For-TI-83-And-TI-84/3000-20415_4-75448635.html?part=dl-&subj=dl&tag=button